OpenStudy (anonymous):

How do I integrate (e^3^x + 5e^-3^x)?

OpenStudy (anonymous):



geerky42 (geerky42):

HINT: \[\Large \int\limits e^{ax} dx = \dfrac{e^{ax}}{a}\]

geerky42 (geerky42):

Think about it, integral is reverse of derivative. if you take a derivative of \(\large \dfrac{e^{ax}}{a}\), you will go back to \(\large e^{ax}\)

geerky42 (geerky42):

Do that helps?

OpenStudy (anonymous):

Oh yeah I see

OpenStudy (anonymous):

forgot that e^kx integrates to 1/k e^kx

geerky42 (geerky42):

yep, so you are good? Can you integrate this?

OpenStudy (anonymous):

yeah, thanks again

geerky42 (geerky42):

welcome!